Define and Describe Quantum Numbers Question Which set of quantum numbers is invalid? Select the cyrect answer below: 1 0 n=2,l=0,m = 0, m, =-흘 n = 3,1= 1,my=0, m, 2 MORE I
media%2F2f8%2F2f8fc2bd-ff14-44c7-b6c7-da
Answer #1

numbers addresses the en elections Quantum atom, o There are principle f size and four Quantum numbers : Quantum number (n) a

3) n=3, l=1, me=0 m. = 2 for n=3 l = (n-1) = (3-1) = 2,1,o there e=1 so, for e=1 me = +10-1 mg = + or - so, this set of Quant